OpenStudy (anonymous):

Let f(x)=3x^2+4x-2 g(x)=x^2-3 Find: f(x+h)

OpenStudy (anonymous):

substitute (x+h) everywhere there is an x in f(x)

OpenStudy (anonymous):

what's the g(x)=x^2-3 doing?

OpenStudy (anonymous):

might be a setup for several problems

OpenStudy (anonymous):

yeah, maybe

OpenStudy (akshay_budhkar):

Y is d person not reponding?

OpenStudy (anonymous):

f(x+h) =3(x+h)^2 + 4(x+h) -2 = 3(x^2+2xh+h^2) +4x +4h -2 = 3x^2 +6xh +3h^2 +4x +4h -2

OpenStudy (anonymous):

thanks so much!!

OpenStudy (anonymous):

You are very welcome! - Good luck